OpenStudy (anonymous):

In trigonometry, cofunctions (sine and cosine, cosecant and secant, and tangent and cotangent) of __________________________ angles are congruent.

OpenStudy (anonymous):

A) complementary B) right C) straight D) supplementary

zepdrix (zepdrix):

Lemme give you an example that might reveal the answer to you,\[\Large\sf \sin(60^o)=\cos(30^o)\] 30 and 60 are what? Complimentary? Supplementary? ..?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ill try

OpenStudy (ipwnbunnies):

You'll try after zepdrix attempts to help you? Basically gives you the answer?

OpenStudy (anonymous):

complementary they make 90 degrees for same value

OpenStudy (anonymous):

i mean ill try to know why is that answer

OpenStudy (anonymous):

oh kk

OpenStudy (anonymous):



OpenStudy (anonymous):

simple minded way to see it \[\cos(\beta)=\frac{a}{h}=\sin(\alpha)\]

zepdrix (zepdrix):

So someone spilled the beans already. :) 30 and 60 and `complimentary` angles. I was trying to show that the cofunctions(sine/cosine) of `complimentary` angles are equal to one another. Able to make sense of this?